OpenStudy (anonymous):

Find the product of 4i and 6i.

OpenStudy (anonymous):

\[4\times 6=24\\ i\times i=-1\]

OpenStudy (anonymous):

-24

Nnesha (nnesha):

\[i \times i = i ^{2}\] \[i ^{2} \]

OpenStudy (anonymous):

i^2=-1 -1*24=-24

Nnesha (nnesha):

\[i \times i = i ^{2}\] \[i ^{2} = -1\]

Nnesha (nnesha):

 now multiply